Band Theory of Correlated Electron Systems

Walter Temmerman

Daresbury Laboratory, Daresbury, Warrington WA4 4AD, U.K.

(W.M.Temmerman@dl.ac.uk)

A review will be given of the self-interaction corrected local spin density (SIC-LSD) method to describe the electronic and magnetic properties of strongly correlated d and f electron systems. Applications of the method for the determination of the valency of rare earths and actinides will be presented. Recent results on the metal/insulator transition in YBCO, half metallicity of Sr_{2}FeMoO_{6} and orbital ordering in LaMnO_{3} will be discussed.